# Flat band effects on the ground-state BCS-BEC crossover in atomic Fermi gases in a quasi-two-dimensional Lieb lattice

### v2

@@ Abstract
The ground-state superfluid behavior of ultracold atomic Fermi gases with an on-site attractive interaction in a quasi-two-dimensional Lieb lattice is studied using BCS mean-field theory, within the context of BCS-BEC crossover. We find that the flat band leads to nontrivial exotic effects. As the Fermi level enters the flat band, both the pairing gap and the in-plane superfluid density exhibit an unusual power law as a function of interaction, with strongly enhanced quantum geometric effects, in addition to a dramatic increase of compressibility as the interaction approaches the BCS limit. As the Fermi level crosses the van Hove singularities, the character of pairing changes from particle-like to hole-like or vice versa. We present the computed phase diagram, in which a pair density wave state emerges at high densities with relatively strong interaction strength.
@@
Ultracold atomic Fermi gases in optical lattices have been of great interest over the past two decades, due to their multiple adjustable parameters, which renders them an ideal platform for quantum simulation and quantum engineering and thus have enormous potential for adding to our understanding of difficult condensed matter problems and exploring new quantum physics [1, 2, 3, 4, 5]. These adjustable parameters include interaction strength, lattice potential well depth, temperature, dimensionality, population imbalance, and lattice geometry, etc [6, 7, 8, 9, 10]. In particular, crossover from a BCS type of superfluidity to Bose–Einstein condensation (BEC) of fermion pairs in an attractive Hubbard model can be realized in an optical lattice using atomic Fermi gases in an optical lattice [11]. Such a BCS-BEC crossover has been realized by tuning the effective interaction strength through a Feshbach resonance in trapped atomic Fermi gases [12]. Such crossover studies can help to elucidate the underlying physics of the widespread pseudogap phenomenon in cuprate superconductors [13], which is of central importance in understanding the mechanism of high-temperature superconductivity [14].
@@
Recently, models with a flat band have attracted great interest, because of the high density of states of the flat band and possible quantum geometric effect associated with multiband of such systems. It has been reported that flat band and quantum geometric effect may enhance the superconducting transition temperature with an on-site attractive interaction and may give rise to quantum Hall states with a nonzero Chern number [15, 16, 17, 18, 19, 20]. Flat bands have been studied in bipartite lattices, e.g., Lieb lattice, perovskite lattice, magic-angle graphene moiré lattices, as well as kagome and honeycomb lattices [21, 22, 15, 23, 24].
@@
Particularly, Lieb lattices have been realized in optical lattices of ultracold atoms [25, 26]. With line centers on a square, a Lieb lattice contains a central flat band, which touches an upper and a lower band at the Dirac points at the Brillouin zone corners [27]. There have been a number of theoretical and experimental studies on Lieb lattices, including lattice preparation [25, 26, 28] and associated aspects including Chern semimetals with three spinless fermion species [29], ferromagnetic and antiferromagnetic states in a repulsive Hubbard model at half filling [30, 31, 32], strain-induced superconductor-insulator transition [33], and competition between pairing and charge density wave at half filling using determinant quantum Monte Carlo [34]. Moreover, quantum spin Hall effect has also been investigated in topological Lieb lattices, when the next-nearest-neighbor hopping associated with spin-orbit-coupling [35, 36, 19] or a Dzyaloshiskii-Moriya term is introduced [37]. Due to the rich physics in the presence of a Lieb lattice, it is important to investigate the superfluidity and pairing phenomenon of ultracold Fermi gases in a Lieb lattice, in order to uncover possible exotic and interesting new quantum phenomena in an attractive Hubbard model in the presence of a flat band.
@@
In this Letter, we investigate the flat band effects on the ground-state superfluid behaviors of ultracold atomic Fermi gases in a quasi-two-dimensional Lieb lattice within a simple attractive Hubbard model, with the nearest neighbor hopping only, which leads to a zero Chern number for the flat band [38]. Among the three bands, we find that the flat band, as well as the van Hove singularities (VHS), strongly affects the superfluid behaviors, leading to many exotic phenomena. When the Fermi level enters the flat band, the fermion pairing gap changes from an exponential dependence into an unusual power law, as a function of the interaction strength, with an enormous compressibility $\kappa$ in the BCS limit. At the same time, the superfluid density also exhibits a power law behavior, in contrast to being a constant in 3D free space. For a lower number density slightly above 1 fermion per unit cell, the fermion chemical potential $\mu$ exhibits a nonmonotonic behavior as a function of the interaction strength in the weak interaction region, and reaches a maximum when it crosses the VHS in the lower band, which signals a change of character of the pairing from particle-like at strong interactions to hole-like at weak interactions. At the same time, the compressibility $\kappa$ is raised up in the weak coupling region as the Fermi level gets closer to the VHS in the lower band. Furthermore, in the BEC regime, $\mu$ asymptotically shows a linear dependence on the interaction strength $|g|$, leading to a density-independent $\kappa=-2/g$, with the pairing gap $\Delta\sim|\mu|\sim|g|$, which is qualitatively similar to the 3D lattice case [39]. Moreover, the ground-state phase diagram reveals that a pair density wave (PDW) ground state emerges at intermediate pairing strength for relatively large densities, as a consequence of strong inter-pair repulsive interactions and relatively large pair size at intermediate pairing strength, which is also found in dipolar Fermi gases [40] and Fermi gases in 2D optical lattice with 1D continuum dimension [41, 42] within the pairing fluctuation theory.
The non-interacting Hamiltonian of a Lieb lattice under the orbital representation is given by
$H_{0}=\sum_{\mathbf{k}\sigma}\hat{c}_{\mathbf{k}\sigma}^{{\dagger}}\hat{H}_{\mathbf{k}}\hat{c}_{\mathbf{k}\sigma}$
@@
where $\hat{c}^{T}_{\mathbf{k}\sigma}=[C_{A\mathbf{k}\sigma},C_{B\mathbf{k}\sigma},C_{C\mathbf{k}\sigma}]$. The symbol A represents the lattice site in a standard square lattice with B and C located on every side of the square, as shown in Fig. (1A) of Ref. [25]. This leads to the Hamiltonian in momentum space for free fermions
@@
where $a_{k}=2t[1-\cos(k_{x}/2)]$ and $b_{k}=2t[1-\cos(k_{y}/2)]$ represent the hopping in the $x$ and $y$ directions, respectively, and $d_{k}=2t_{z}(1-\cos k_{z})-\mu$ the dispersion in the out-of-plane $\hat{z}$ direction, with $t$ and $t_{z}$ being the in-plane and out-of-plane hopping integral, respectively. We take $t_{z}/t=0.01$ for the quasi-two dimensionality, set the lattice constant $a=1$, and take the bottom of the lower energy band as the zero point of $\mu$. Diagonalizing $\hat{H}_{\mathbf{k}}$ leads to three bands with dispersions $\xi_{\mathbf{k}}^{\alpha}=\alpha\sqrt{2}t\sqrt{2+\cos k_{x}+\cos k_{y}}+2\sqrt{2}t+2t_{z}(1-\cos k_{z})-\mu$, where $\alpha=\pm,0$ denotes the upper, the lower and the flat band, respectively. This yields $H_{0}$ in the band representation, $H_{0}=\sum_{\mathbf{k}\alpha\sigma}\xi^{\alpha}_{\mathbf{k}}c_{\mathbf{k}\alpha\sigma}^{{\dagger}}c_{\mathbf{k}\alpha\sigma}$ , where $c_{\mathbf{k}\alpha\sigma}$ is the annihilation operator in band $\alpha$.
@@
where $\alpha,\beta=\pm,0$. Since the nearest-neighbor hopping hybridizes different orbitals, we find it reasonable to assume the matrix elements of the on-site pairing interaction to be uniform, with $U_{\mathbf{k}\mathbf{k}^{{}^{\prime}}\alpha\beta}=g<0$. Using the mean field approximation in the band representation, this leads to a uniform order parameter $\Delta_{\alpha}=\Delta$ [43]. We emphasize that, despite the short interaction range, fermion hopping enables pairing between different lattice sites across a large distance.
@@
$G^{-1}(K)$ $=$ $G_{0}^{-1}(K)-\Delta^{2}\tilde{G}_{0}(K)\,,$
respectively, where $\tilde{G}_{0}(K)=-G_{0}^{T}(-K)$ with four momentum $K\equiv(\omega,\mathbf{k})$, and $\theta(x)$ is the Heaviside step function. Thus one has the self energy $\Sigma(K)=\Delta^{2}\tilde{G}_{0}(K)$.
Under the number constraint $n=2\sum_{K}{\rm Tr}\,G(K)$ with $\sum_{K}\equiv\sum_{\omega}\sum_{\mathbf{k}}$, we have the number equation
@@
where $E_{\mathbf{k}}^{\alpha}=\sqrt{(\xi_{\mathbf{k}}^{\alpha})^{2}+\Delta^{2}}$ is the Bogoliubov quasiparticle dispersion for the $\alpha=\pm,0$ band with an energy gap $\Delta$. And the gap equation is given by[43]
@@
Equations (1) and (2) form a closed set of self-consistent equations, which can be solved for ($\mu$, $\Delta$) as a function of $g$ in the superfluid phase.
Furthermore, the solution for ($\mu$, $\Delta$) from Eqs.(1) and (2) should satisfy the stability condition for the superfluid phase that the energy spectrum of of Cooper pairs should be nonnegative. To this end, we extract the pair dispersion using the fluctuating pair propagator, as given in the pairing fluctuation theory which was previously developed for the pseudogap physics in the cuprates [14], and has been extended to address the BCS-BEC crossover in ultracold atomic Fermi gases [1]. To be compatible [44] with the BCS-Leggett ground state, the pairing susceptibility in the $T$-matrix $t_{\text{pg}}^{-1}(Q)=1/g+\chi(Q)$ is given by $\chi(Q)=-\sum_{K}{\rm Tr}\,[G(K)\tilde{G}_{0}(K\!-Q)]$, with $Q\equiv(\Omega,\mathbf{\mathbf{q}})$, which leads to $t_{\text{pg}}^{-1}(\Omega,\mathbf{q})\approx a_{0}(\Omega-\Omega_{\mathbf{q}})$ in the superfluid phase, with the pair dispersion $\Omega_{\mathbf{q}}=2B(2-\cos q_{x}-\cos q_{y})+2B_{z}(1-\cos q_{z})$. Here $B$ and $B_{z}$ correspond to the effective pair hopping integral in the $xy$ plane and in the $z$ direction, respectively. The expressions for the coefficients $a_{0}$, $B$ and $B_{z}$ can be readily derived during the Taylor expansion. The non-negativeness of the pair dispersion requires that the pairing correlation length (squared) $\xi^{2}=a_{0}B$ and $\xi_{z}^{2}=a_{0}B_{z}$ be positive.
Superfluid density is an important transport property, which is correlated with condensed pairs; its dependence on the interaction strength often reflects the pairing symmetry. It takes the average of the inverse band mass in the presence of a lattice, in contrast with the 3D continuum case where it is always given by $n/m$ at zero $T$ in BCS theory. Moreover, the in-plane component of the superfluid density can be divided into a conventional and a geometric part, due to the existence of the flat band, where the geometric term is associated with the interband matrix elements of the current operator, which is proportional to a quantum metric tensor.
The expressions for superfluid density are derived using the linear response theory within the BCS framework [45, 46], which is applied to the multi-band system [47] [^1] Note that the $(n_{s}/m)$ tensor has to be calculated as a whole, not $n_{s}$ and $m$ separately.. Similar to Ref. [47], the in-plane superfluid density $(n_{s}/m)_{\parallel}$ contains a conventional term $(n_{s}/m)^{\rm conv}_{\parallel}$ and a geometric term $(n_{s}/m)_{\parallel}^{\rm geom}$, i.e.,
@@
Compressibility $\kappa$ is one of the most important physical quantities in thermodynamic physics, which must be positive to maintain mechanical stability. For the ground-state compressibility [49], one has
$\begin{split}\kappa&=\frac{\partial n}{\partial\mu}=\left(\frac{\partial n}{\partial\mu}\right)_{\Delta}+\left(\frac{\partial n}{\partial\Delta}\right)_{\mu}\frac{\partial\Delta}{\partial\mu}\\ &=\sum_{k\alpha}\frac{\Delta^{2}}{(E_{\mathbf{k}}^{\alpha})^{3}}+\frac{\left[\sum_{k\alpha}{\xi_{\mathbf{k}}^{\alpha}}/{(E_{\mathbf{k}}^{\alpha})^{3}}\right]^{2}}{\sum_{k\alpha}{1}/{(E_{\mathbf{k}}^{\alpha})^{3}}}\,.\end{split}$ (7)
@@
$\mu$ $=$ $\frac{(3-n)g}{2}+2\sqrt{2}t\,,$ (8)
$\Delta$ $=$ $\sqrt{\frac{9}{4}g^{2}-\mu^{2}}\,,$ (9)

### v3

@@ Abstract
The ground-state superfluid behavior of ultracold atomic Fermi gases with a short-range attractive interaction in a quasi-two-dimensional Lieb lattice is studied using BCS mean-field theory, within the context of BCS-BEC crossover. We find that the flat band leads to nontrivial exotic effects. As the Fermi level enters the flat band, both the pairing gap and the in-plane superfluid density exhibit an unusual power law as a function of interaction, with strongly enhanced quantum geometric effects, in addition to a dramatic increase of compressibility as the interaction approaches the BCS limit. As the Fermi level crosses the van Hove singularities, the character of pairing changes from particle-like to hole-like or vice versa. We present the computed phase diagram, in which a pair density wave state emerges at high densities with relatively strong interaction strength.
I Introduction
@@
Ultracold atomic Fermi gases in optical lattices have become an ideal platform for quantum simulation and quantum engineering and thus have enormous potential for exploring difficult condensed matter problems and new quantum physics [1, 2, 3, 4, 5], due to their multiple adjustable parameters, including interaction strength, lattice depth, dimensionality, population imbalance, and lattice geometry, etc [6, 7, 8, 9, 10]. In particular, crossover from a BCS type of superfluidity to Bose–Einstein condensation (BEC) of fermion pairs in an attractive Hubbard model can be realized in an optical lattice using atomic Fermi gases in an optical lattice [11]. Such a BCS-BEC crossover has been realized by tuning the effective interaction strength through a Feshbach resonance in trapped atomic Fermi gases [12]. Such studies can help to elucidate the underlying physics of the widespread pseudogap phenomena in cuprate superconductors [13], which is of central importance in understanding the mechanism of high-temperature superconductivity [14].
@@
Recently, models with a flat band have attracted great interest, because of the associated high density of states (DOS) and possible quantum geometric effect associated with multiband of such systems. It has been reported that flat band and quantum geometric effect may enhance the superconducting transition temperature with an on-site attractive interaction and may give rise to quantum Hall states with a nonzero Chern number [15, 16, 17, 18, 19, 20]. Flat bands have been studied in bipartite lattices, e.g., Lieb lattice, magic-angle graphene moiré lattices, as well as perovskite, kagome and honeycomb lattices [21, 22, 15, 23, 24].
@@
Particularly, Lieb lattices have been realized in optical lattices of ultracold atoms [25, 26]. With line centers on a square, a Lieb lattice contains a central flat band, which touches an upper and a lower band at the Dirac points at the Brillouin zone corners [27]. There have been a number of theoretical and experimental studies on Lieb lattices, including lattice preparation [25, 26, 28] and associated aspects such as Chern semimetals with three spinless fermion species [29], ferromagnetic and antiferromagnetic states in a repulsive Hubbard model at half filling [30, 31, 32], strain-induced superconductor-insulator transition [33], competition between pairing and charge density wave at half filling using determinant quantum Monte Carlo [34], as well as topologically nontrivial quantum spin Hall effect with extra interactions [35, 36, 19, 37]. Such rich physics warrants the investigation of the important superfluidity and pairing phenomenon of ultracold Fermi gases in a Lieb lattice, in order to uncover possible exotic and interesting new quantum phenomena in the presence of a flat band.
@@
In this Letter, we investigate the flat band effects on the ground-state superfluid behaviors of ultracold atomic Fermi gases in a quasi-two-dimensional Lieb lattice with a short-range attractive interaction, with the nearest neighbor hopping only, which leads to a zero Chern number [38]. We find that the flat band, as well as the van Hove singularities (VHS), leads to exotic phenomena in the superfluid behavior. When the Fermi level enters the flat band, the fermion pairing gap $\Delta$ changes from an exponential dependence into an unusual power law, as a function of the interaction strength $|g|$, with an enormous compressibility $\kappa$ in the BCS limit. The superfluid density also exhibits a power law behavior, in contrast to being a constant in 3D free space. For a lower number density $n$ slightly above 1 fermion per unit cell, the fermion chemical potential $\mu$ varies nonmonotonically with the interaction strength in the weak interaction region, and reaches a maximum when it crosses the VHS, signaling a change of pairing character from particle-like at strong interactions to hole-like at weak interactions. At the same time, $\kappa$ rises in the weak coupling region as the Fermi level gets closer to the VHS. In the BEC regime, $\mu$ shows an asymptotic linear dependence on the interaction $g$, leading to a density-independent $\kappa=-2/g$, with the pairing gap $\Delta\sim|\mu|\sim|g|$, which is qualitatively similar to the 3D lattice case [39]. Finally, the phase diagram reveals that a pair density wave (PDW) ground state emerges at intermediate pairing strength for relatively large densities, due to strong inter-pair repulsive interactions and relatively large pair size at intermediate pairing strength, which is also found in dipolar Fermi gases [40] and Fermi gases in 2D optical lattices with one continuum dimension [41, 42] within the pairing fluctuation theory.
II Theoretical Formalism
Figure 1: Structure of a quasi-2D Lieb lattice. Site A represents the site in a simple square lattice, while sites B and C sit at the line centers of the square in each 2D plane. The solid lines represent nearest-neighbor hopping integral $t$ between site A and sites B and C. The dashed lines denote the inter-layer coupling $t_{z}$.
The non-interacting Hamiltonian of a Lieb lattice can be written as
$H_{0}=\sum_{\mathbf{k}\sigma}\hat{c}_{\mathbf{k}\sigma}^{{\dagger}}\hat{H}_{\mathbf{k}}\hat{c}_{\mathbf{k}\sigma},$
@@ II Theoretical Formalism
where $\hat{c}_{\mathbf{k}\sigma}=[C_{A\mathbf{k}\sigma},C_{B\mathbf{k}\sigma},C_{C\mathbf{k}\sigma}]^{T}$. The subscript A denotes the site in a standard square lattice with B and C located on the side center of the square, as shown in Fig. 1. This leads to the Hamiltonian in momentum space for free fermions
@@
where $a_{k}=2t[1-\cos(k_{x}/2)]$, $b_{k}=2t[1-\cos(k_{y}/2)]$ and $d_{k}=2t_{z}(1-\cos k_{z})-\mu$ represent the hopping in the $x$ and $y$ directions, and the dispersion in the out-of-plane $\hat{z}$ direction, respectively, with $t$ and $t_{z}$ being the in-plane and out-of-plane hopping integral, respectively. We take $t_{z}/t=0.01$ for the quasi-two dimensionality, set the lattice constant $a=1$, and measure energy relative to the bottom of the lower energy band. Diagonalizing $\hat{H}_{\mathbf{k}}$ leads to three bands with dispersions $\xi_{\mathbf{k}}^{\alpha}=\alpha\sqrt{2}t\sqrt{2+\cos k_{x}+\cos k_{y}}+2\sqrt{2}t+2t_{z}(1-\cos k_{z})-\mu$, where $\alpha=\pm,0$ denotes the upper, lower and flat band, respectively. This yields in the band representation $H_{0}=\sum_{\mathbf{k}\alpha\sigma}\xi^{\alpha}_{\mathbf{k}}c_{\mathbf{k}\alpha\sigma}^{{\dagger}}c_{\mathbf{k}\alpha\sigma}$, where $c_{\mathbf{k}\alpha\sigma}$ is the annihilation operator in band $\alpha$.
@@
with band indices $\alpha,\beta=\pm,0$. The full Hamiltonian is thus $H=H_{0}+H_{\mathrm{int}}$. We find that with only on-site interactions, the pairing gap often varies strongly from site to site [43], which necessarily leads to a large kinetic energy for the pairing field, and thus may not be energetically favorable in the superfluid state. Since the nearest-neighbor hopping hybridizes different orbitals, and to be compatible with the conventional superconductivity with electron-phonon interaction induced pairing, we find it reasonable to assume a uniform short-range intra- and inter-orbital interaction, $U_{ij}=U$, where $i,j=\{A,B,C\}$ are the orbital indices. Through a unitary transformation, this leads naturally to uniform matrix elements of the interaction in momentum space across all bands, with $U_{\mathbf{k}\mathbf{k}^{{}^{\prime}}\alpha\beta}=g<0$. (See Appendix A for details). This in turn gives rise to a uniform order parameter $\Delta_{\alpha}=\Delta$ in the mean field approximation [44]. We emphasize that, despite the short interaction range, fermion hopping enables pairing between sites across a large distance.
@@
$G^{-1}(K)$ $=$ $G_{0}^{-1}(K)-\Sigma(K),\quad\Sigma(K)=-\Delta^{2}{G}^{T}_{0}(-K)\,,$
respectively, with four momentum $K\equiv(\omega,\mathbf{k})$, and $\theta(x)$ is the step function.
The number constraint $n=2\sum_{K}{\rm Tr}\,G(K)$ leads to the number equation
@@
where $\sum_{K}\equiv\sum_{\omega}\sum_{\mathbf{k}}$, and $E_{\mathbf{k}}^{\alpha}=\sqrt{(\xi_{\mathbf{k}}^{\alpha})^{2}+\Delta^{2}}$ is the Bogoliubov quasiparticle dispersion for band $\alpha=\pm,0$ with energy gap $\Delta$. The gap equation is given by [44]
@@
Equations (1) and (2) form a closed set of self-consistent equations , which can be solved for ($\mu$, $\Delta$) as a function of $g$ in the superfluid phase. The solution for ($\mu$, $\Delta$) in the superfluid phase should satisfy the stability condition that the Cooper pair energy should be nonnegative. To this end, we extract the pair dispersion using the fluctuating pair propagator, as given in the pairing fluctuation theory which was previously developed for the pseudogap physics in the cuprates [14], and extended to address the BCS-BEC crossover in atomic Fermi gases [1]. To be compatible [45] with the BCS-Leggett ground state, the pair susceptibility in the $T$-matrix $t_{\text{pg}}^{-1}(Q)=1/g+\chi(Q)$ is given by $\chi(Q)=\sum_{K}{\rm Tr}\,[G(K){G}^{T}_{0}(Q-K)]$, with $Q\equiv(\Omega,\mathbf{\mathbf{q}})$, which leads to $t_{\text{pg}}^{-1}(\Omega,\mathbf{q})\approx a_{0}(\Omega-\Omega_{\mathbf{q}})$, with pair dispersion $\Omega_{\mathbf{q}}=2B(2-\cos q_{x}-\cos q_{y})+2B_{z}(1-\cos q_{z})$. Here $B$ and $B_{z}$ correspond to the effective pair hopping integral in the $xy$ plane and in the $z$ direction, respectively. The expressions for the coefficients $a_{0}$, $B$ and $B_{z}$ can be readily derived during the Taylor expansion. The non-negativeness of the pair dispersion requires that the pairing correlation length (squared) $\xi^{2}=a_{0}B$ and $\xi_{z}^{2}=a_{0}B_{z}$ be positive.
Superfluid density $(n_{s}/m)$ is an important transport property; its dependence on the interaction often reflects the pairing symmetry. It takes the average of the inverse band mass in a lattice, in contrast with the 3D continuum case where it is always given by $n/m$ at $T=0$ in BCS theory. Moreover, the in-plane component of $(n_{s}/m)$ can be divided into a conventional and a geometric part, due to the presence of the flat band, where the geometric term is associated with the interband contributions, which are proportional to a quantum metric tensor.
The expressions for the superfluid density $(n_{s}/m)$ are derived using the linear response theory within the BCS framework [46, 47], which is applied to the multi-band system [48] [^1] Note that the $(n_{s}/m)$ tensor has to be calculated as a whole, not $n_{s}$ and $m$ separately.. Similar to Ref. [48], the in-plane superfluid density $(n_{s}/m)_{\parallel}$ contains a conventional term $(n_{s}/m)^{\rm conv}_{\parallel}$ and a geometric term $(n_{s}/m)_{\parallel}^{\rm geom}$, i.e.,
@@
Note that, without using the mean-field approximated Hamiltonian, we do not find the extra contributions to $n_{s}/m$ associated with the derivative of $\Delta$ with respect to the magnetic field vector potential, unlike Ref. [50].
Compressibility $\kappa$ is an important quantity in thermodynamics, which must be positive to maintain mechanical stability. For the ground state [51], we obtain
$\kappa$ $=\frac{\partial n}{\partial\mu}=\left(\frac{\partial n}{\partial\mu}\right)_{\Delta}+\left(\frac{\partial n}{\partial\Delta}\right)_{\mu}\frac{\partial\Delta}{\partial\mu}$
$=\sum_{k\alpha}\frac{\Delta^{2}}{(E_{\mathbf{k}}^{\alpha})^{3}}+\frac{\left[\sum_{k\alpha}{\xi_{\mathbf{k}}^{\alpha}}/{(E_{\mathbf{k}}^{\alpha})^{3}}\right]^{2}}{\sum_{k\alpha}{1}/{(E_{\mathbf{k}}^{\alpha})^{3}}}\,.$ (7)
@@
$\mu=\frac{(3-n)g}{2}+2\sqrt{2}t\,,\quad\Delta=\sqrt{\frac{9}{4}g^{2}-\mu^{2}}\,,$ (8)